OpenStudy (anonymous):

Look at the right triangle ABC. A student made the following chart to prove that AB2 + BC2 = AC2.

OpenStudy (anonymous):

1. Triangle ABC is similar to triangle BDC 1. Angle ABC = Angle BDC and Angle BCA = Angle BCD 2. BC2 = AC × DC 2. BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC 3. Triangle ABC is similar to triangle ABD 3. Angle ABC = Angle BAD and Angle BAC = Angle ABD 4. AB2 = AC × AD 4. AB ÷ AD = AC ÷ AB because triangle ABC is similar to triangle ABD 5. AB2 + BC2 = AC × AD + AC × DC = AC (AD + DC) 5. Adding Statement 1 and Statement 2 6. AB2 + BC2 = AC2 6. AD + DC = AC

OpenStudy (anonymous):

What is the flaw in the student’s proof? Justification 4 should be “AB ÷ AD = AB ÷ AC because triangle ABC is similar to triangle ABD”. Justification 1 should be ““Angle ABC = Angle BCD and Angle BCA = Angle DBC”. Justification 2 should be “BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC”. Justification 3 should be “Angle ABC = Angle ADB and Angle BAC = Angle BAD”.

OpenStudy (anonymous):

lets party!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

Justification 2 should be “BC ÷ DC = AC ÷ BC because triangle ABC is similar to triangle BDC”.

OpenStudy (anonymous):

I think justification 2 looked funny

OpenStudy (anonymous):

what do you guys think?

OpenStudy (anonymous):

oh crap

OpenStudy (anonymous):

thats the wrong answer

OpenStudy (anonymous):

Is there a drawing or can you put a picture of the chart?

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

lol the file name...

OpenStudy (anonymous):

pizza hut!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Well, sorry. I am not familiar with this kind of proof...even with the cool pizza name...maybe next time?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I'm calling out the big guns now...

OpenStudy (anonymous):

looks like we may need a scorpion vs a wraith

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

don't worry, cal, I'm calling Zepp

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Zepp

OpenStudy (anonymous):

could you help me with this type of problem?

OpenStudy (zepp):

I did that proof last week .-. forgot how xD lemme see

OpenStudy (anonymous):

k

OpenStudy (anonymous):

There's actually a guy named lolJDlol with a creeper profile pic

OpenStudy (anonymous):

Its boss

OpenStudy (anonymous):

and creepy at the same time

OpenStudy (zepp):

Basically, this student is making a proof on the pythagorean's theorem by using similar triangles.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

.so..

OpenStudy (anonymous):

Justification 4 should be “AB ÷ AD = AB ÷ AC because triangle ABC is similar to triangle ABD”.

OpenStudy (anonymous):

>thgirir

OpenStudy (anonymous):

right?

OpenStudy (zepp):

3. Angle ABC = Angle BAD and Angle BAC = Angle ABD 4. AB2 = AC × AD This is wrong.

OpenStudy (zepp):

Angle ABC =/= BAD

OpenStudy (anonymous):

so justification 4 is wrong

OpenStudy (anonymous):

k

OpenStudy (zepp):

:)

OpenStudy (anonymous):

k, going to take the test

OpenStudy (anonymous):

hope I do well

OpenStudy (zepp):

Have fun and do well! :)